# Leave No One Behind: Online Self-Supervised Self-Distillation for Sequential Recommendation

### v1

@@ Abstract.
Sequential recommendation methods play a pivotal role in modern recommendation systems. A key challenge lies in accurately modeling user preferences in the face of data sparsity. To tackle this challenge, recent methods leverage contrastive learning (CL) to derive self-supervision signals by maximizing the mutual information of two augmented views of the original user behavior sequence. Despite their effectiveness, CL-based methods encounter a limitation in fully exploiting self-supervision signals for users with limited behavior data, as users with extensive behaviors naturally offer more information. To address this problem, we introduce a novel learning paradigm, named Online Self-Supervised Self-distillation for Sequential Recommendation ($S^{4}$ec), effectively bridging the gap between self-supervised learning and self-distillation methods. Specifically, we employ online clustering to proficiently group users by their distinct latent intents. Additionally, an adversarial learning strategy is utilized to ensure that the clustering procedure is not affected by the behavior length factor. Subsequently, we employ self-distillation to facilitate the transfer of knowledge from users with extensive behaviors (teachers) to users with limited behaviors (students). Experiments conducted on four real-world datasets validate the effectiveness of the proposed method [^1] Code is available at https://github.com/xjaw/S4Rec.
Sequential Recommendation; Multi-Intention Modeling; Self-Supervised Learning; Long-Tail Learning
journalyear: 2024copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2024; May 13–17, 2024; Singapore, Singapore.booktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporeisbn: 978-1-4503-XXXX-X/18/06doi: XXXXXXX.XXXXXXXccs: Computing methodologies Learning paradigmsccs: Computing methodologies Neural networksccs: Information systems Recommender systems
@@ 2.1. Self-Supervised Learning for Recommendation
[…]
@@ 2.2. Intention Learning for Recommendation
Many approaches have been proposed to study users’ intents behind each user’s behavior for improving recommendations (Ma et al., 2020; Chen et al., 2020; Tan et al., 2021). DSSRec (Ma et al., 2020) introduces a sequence2sequence training strategy to capture extra supervision information. An intent variable is employed to extract mutual information between an individual user’s past and future interaction sequences. ICLRec (Chen et al., 2022) learns users’ intent distribution from unlabeled user behavior sequences and optimize SR models with contrastive learning by considering the obtained intents. ISRec (Li et al., 2023b) extracts intentions of target user from sequential contexts, then takes complex intent transition into account through the message-passing mechanism on intention graphs.
@@ 4. METHODOLOGY
In this section, we discuss the details of our proposed $S^{4}$ec. The overall framework is illustrated in Figure 2.
Figure 2. The overall framework of $S^{4}$ec. It augments original behavior sequences as contrastive views and employs online clustering to proficiently group users by their distinct latent intents. Subsequently, $S^{4}$ec conducts cluster-aware self-distillation to transfer knowledge from corresponding intents (teachers) that contain extensive behavior information to users with limited behaviors (students). Additionally, an adversarial learning strategy ensures that the the clustering procedure is not affected by the head-tail factor. The right part describes the online clustering process based on the Sinkhorn-Knopp algorithm.
@@ 4.1. Clustering On The Fly
Then the soft cluster assignment $\mathbf{p}_{u}$ of $u$ can be calculated as :
(3) $\centering\begin{split}\mathbf{p}_{u}^{k}=\frac{\exp(\mathbf{z}_{u}\bm{\mu}_{k}^{\top}/\tau_{1})}{\sum_{k^{\prime}}\exp(\mathbf{z}_{u}\bm{\mu}_{k^{\prime}}^{\top}/\tau_{1})},\end{split}\@add@centering$
where $\bm{\mu}_{k}$ is the $k$-th row of $\bm{\mu}\in\mathbb{R}^{K\times d}$, which represents $K$ trainable prototypes, i.e. intent representations. $\tau_{1}$ is a temperature parameter.
We then further refine the cluster assignment with the help of an auxiliary target distribution $\mathbf{q}_{u}$, obtained by mapping $\mathbf{z}_{u}$ to $\bm{\mu}$. The objective is trained by matching the soft assignment to the target distribution, specifically defined as a cross entropy loss between the soft assignment $\mathbf{p}_{u}$ and the auxiliary target $\mathbf{q}_{u}$, i.e.,
(4) $\centering\begin{split}\mathcal{L}_{Clust}(\mathbf{z}_{u},\mathbf{q}_{u})=-\sum_{k}\mathbf{q}_{u}^{k}\log\mathbf{p}_{u}^{k}.\end{split}\@add@centering$
The objective function is jointly minimized with respect to the prototypes $\bm{\mu}$ and the parameters $\theta$ of the sequence encoder $f_{\theta}$ used to produce the sequence representation $\mathbf{z}_{u}$.
We now introduce the method to obtain the auxiliary target and update the prototypes. To make our proposal cluster online, we iteratively compute the auxiliary target using only the sequence features within a batch. We utilize the prototypes $\bm{\mu}$ to compute the auxiliary target and enforce all the instances in a batch equally partitioned by the prototypes as much as possible. This equipartition constraint ensures that the auxiliary targets for different sequences in a batch are distinct, thus preventing the trivial solution where every sequence has the same auxiliary target (Caron et al., 2020).
Given $B$ embedding vectors $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{B}]$ in a mini-batch, we are interested in mapping them to the prototypes $\bm{\mu}=[\bm{\mu}_{1},...,\bm{\mu}_{K}]$. We denote this mapping or target distribution as $\mathbf{Q}=[\mathbf{q}_{1},...,\mathbf{q}_{B}]$, and optimize $\mathbf{Q}$ to maximize the similarity between sequence embeddings and prototypes:
(5) $\centering\begin{split}\mathop{\max}_{\mathbf{Q}\in\mathcal{Q}}Tr(\mathbf{Q}\bm{\mu}\mathbf{Z}^{\top})+\epsilon H(\mathbf{Q}),\\ H(\mathbf{Q})=-\sum_{ij}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij},\end{split}\@add@centering$
where $H$ is a entropy function and $\epsilon$ is a parameter adjusting the smoothness of the auxiliary target.
We adopt the solution in (Asano et al., 2020; Caron et al., 2020) that achieves an equal partition by modelling the $\mathbf{Q}$ to belong to the transportation polytope within mini-batch:
(6) $\centering\begin{split}\mathbf{Q}=\{\mathbf{Q}\in\mathbb{R}_{+}^{B\times K}|\mathbf{Q}\mathbf{1}_{K}=\frac{1}{B}\mathbf{1}_{B},\mathbf{Q}^{\top}\mathbf{1}_{B}=\frac{1}{K}\mathbf{1}_{K}\},\end{split}\@add@centering$
where $\mathbf{1}_{B}$ denotes the vector of ones in dimension of batch size $B$. These constraints enforce that on average each prototype is selected at least $\frac{B}{K}$ times within a mini-batch.
One solution $\mathbf{Q}^{*}$ of Eq. (5) over the set $\mathcal{Q}$ takes the form of a normalized exponential matrix and is as following (Cuturi, 2013):
(7) $\centering\begin{split}\mathbf{Q}^{*}=diag(\mathbf{m})\exp(\frac{\mathbf{Z}\bm{\mu}^{\top}}{\epsilon})diag(\mathbf{v}).\end{split}\@add@centering$
where $\mathbf{m}\in\mathbb{R}^{B}$ and $\mathbf{v}\in\mathbb{R}^{K}$ are re-normalized vectors that are computed using the iterative Sinkhorn-Knopp algorithm (Cuturi, 2013), which requires only a small number of matrix multiplications.
@@ 4.2. Cluster-aware Self-distillation
4.2.1. Sequence-level Contrastive Learning
Users’ sequential behaviors naturally present extensive information for obtaining self-supervised signals. Given a function set $\mathcal{G}$ of several data augmented operators, such as mask, crop, reorder and insert (Li et al., 2023a) and an user’s sequence $\mathcal{S}_{u}$, we can create two augmented views as:
(8) $\centering\begin{split}\tilde{\mathcal{S}}_{u}^{1}&=g_{u}^{1}(\mathcal{S}_{u}),\tilde{\mathcal{S}}_{u}^{2}=g_{u}^{2}(\mathcal{S}_{u}),g_{u}^{1},g_{u}^{2}\in\mathcal{G},\end{split}\@add@centering$
where $g_{u}^{1}$ and $g_{u}^{2}$ are augmentation functions sampled from $\mathcal{G}$ to produce different views for $\mathcal{S}_{u}$. Generally, views captured from the same sequence are treated as positive pairs, while those of any other sequences are considered as negative pairs. Furthermore, the contrastive views are mapping into representations $\mathbf{z}_{u}^{1}$ and $\mathbf{z}_{u}^{2}$ by the sequence encoder $f_{\theta}$. After that, we can maximize the mutual information to provide self-supervised signals to improve recommendation performance:
(9) $\centering\begin{split}\mathcal{L}_{SCL}=\mathcal{L}_{SCL}(\mathbf{z}_{u}^{1},\mathbf{z}_{u}^{2})+\mathcal{L}_{SCL}(\mathbf{z}_{u}^{2},\mathbf{z}_{u}^{1}),\\ \mathcal{L}_{SCL}(\mathbf{z}^{1}_{u},\mathbf{z}^{2}_{u})=-\log\frac{\exp(\mathbf{z}^{1}_{u}\mathbf{z}^{2\top}_{u}/\tau_{2})}{\sum_{n\neq u}\exp(\mathbf{z}^{1}_{u}\mathbf{z}^{\top}_{n}/\tau_{2})},\end{split}\@add@centering$
where $\mathbf{z}_{n}$ are negative views’ representations of sequence $\mathcal{S}_{u}$ and $\tau_{2}$ is the temperature parameter tuning the strength of penalties on the hard negative samples (Wang and Liu, 2021).
The classical InfoNCE (van den Oord et al., 2018) loss Eq. (9) only utilized the supervision information of the sample itself. Now that we have gained users’ intent, we consider further pushing users’ embedding away from the cluster they do not belong to by adding extra negative intent information. Thus $\mathcal{L}_{SCL}(\cdot,\cdot)$ can be rewritten as:
(10) $\centering\begin{split}&\mathcal{L}_{SCL}(\mathbf{z}^{1}_{u},\mathbf{z}^{2}_{u})=\\ &-\log\frac{\exp(\mathbf{z}^{1}_{u}\mathbf{z}^{2\top}_{u}/\tau_{2})}{\sum_{n\neq u}\exp(\mathbf{z}^{1}_{u}\mathbf{z}^{\top}_{n}/\tau_{2})+\sum_{k^{\prime}\neq h(u)}\exp(\mathbf{z}^{1}_{u}\bm{\mu}^{\top}_{k^{\prime}}/\tau_{2})},\end{split}\@add@centering$
where $h(u)$ is a function maps $u$ into the index of assigned intent.
4.2.2. Cluster-level Distillation Module
Considering the sparsity of most users’ sequences, the capability of supervisory signals within the instance scope is limited. Knowledge Distillation (KD) (Hinton et al., 2015) is a preferred choice for alleviating sparsity problems. However, most KD methods require a large pretrained, and relatively complex teacher model, which cannot be met in all cases.
Confronting the expensive KD process, self-distillation (Yang et al., 2019; Zhang et al., 2019) is proposed to eliminate the requirement of complex teacher model by sharing the same backbone network to serve as teachers and students simultaneously. Hence, the distillation process is significantly simplified. Nevertheless, this work distinguishes teachers and students through distinct sub-network configurations, which is not preferred due to the complexities.
We propose a cluster-aware self-distillation loss that encourages the sequence embedding (student) to be close to the assigned intent distribution (teacher). Remember that the assigned intent of $u$ is denoted as $\bm{\mu}_{h(u)}$. First, we derive the normalized distribution for sequence embedding and corresponding intent embedding:
(11) $\centering\begin{split}\mathbf{e}_{u,t}^{i}=\frac{\exp(\bm{\mu}_{h(u)}^{i}/\tau_{3})}{\sum_{i^{\prime}}\exp(\bm{\mu}_{h(u)}^{i^{\prime}}/\tau_{3})},\\ \mathbf{e}_{u,s}^{i}=\frac{\exp(\mathbf{z}_{u}^{i}/\tau_{3})}{\sum_{i^{\prime}}\exp(\mathbf{z}_{u}^{i^{\prime}}/\tau_{3})},\end{split}\@add@centering$
Here $\tau_{3}$ is a hyper-parameter called distillation temperature, which controls the smoothness of the normalized distribution. Then the cluster-aware self-distillation loss can be directly defined as KL divergence:
(12) $\centering\begin{split}\mathcal{L}_{CKD}&=\mathcal{L}_{CKD}(\bm{\mu}_{h(u)},\mathbf{z}^{1}_{u})+\mathcal{L}_{CKD}(\bm{\mu}_{h(u)},\mathbf{z}^{2}_{u}),\\ &\mathcal{L}_{CKD}(\bm{\mu}_{h(u)},\mathbf{z}^{1}_{u})=\sum_{i}\mathbf{e}_{u,t}^{i}\log\frac{\mathbf{e}_{u,t}^{i}}{\mathbf{e}_{u,s}^{i}},\end{split}\@add@centering$
@@ 4.3. Head-tail Adversarial Learning
To achieve this, we enhance the recommendation model by introducing a classifier that utilizes the learned sequence embeddings as input. The objective is to train the classifier to accurately predict the category of each sequence. Simultaneously, we aim to optimize the encoder to generate sequence embeddings that can effectively deceive the classifier. By jointly training the classifier and the encoder, we strive to eliminate the length information, leading to a more uniform clustering distribution.
In our experiments, we employ a fully connected layer as the classifier for sequence category classification and utilize cross entropy loss to optimize the classifier. The formula is:
(13) $\centering\begin{split}&\hat{\mathbf{c}}=\mathbf{W}\mathbf{z}_{u},\\ \mathcal{L}_{Adv}=-\hat{\mathbf{c}}[c]&+\log(\sum_{i}\exp(\hat{\mathbf{c}}[i])),\end{split}\@add@centering$
where $\hat{\mathbf{c}}$ is output logits of the classifier, and $c$ is the corresponding category of sequence $\mathcal{S}_{u}$. Under the setting of adversarial learning, the object for the sequence category classifier is to minimize $\mathcal{L}_{Adv}$, and the object for the encoder is to minimize $\mathcal{L}_{SR}-\gamma\mathcal{L}_{Adv}$, where $\gamma$ is introduced to balance the main task and the additional adversarial task.
@@
In the context of implementing adversarial learning, one elegant approach is to incorporate a Gradient Reversal Layer (GRL) within the backward propagation process which is initially introduced in DAN (Ganin and Lempitsky, 2015). Since we expect the classifier to minimize $\mathcal{L}_{Adv}$, while forcing the main encoder to maximize $\mathcal{L}_{Adv}$, we insert a GRL layer between the main encoder and the fully connected classifier. During the backpropagation process, the gradients for minimizing the classification loss flow backward through the classifier, and after the GRL, the gradients will be reversed, which further flows to the encoder. That is, we perform gradient descent on the parameters of the classifier while performing gradient ascent on the parameters of the encoder, with respect to $\mathcal{L}_{Adv}$. for other objectives, gradient descent is applied to the encoder. Through this subtle design, we successfully implement the adversarial learning task.
@@ 4.4. Multi-task Training
We adopt a multi-task strategy where the main next-item prediction, the cluster assignment, the cluster-aware self-distillation and the adversarial learning task are jointly optimized. The joint loss is a linear weighted sum calculated as:
(14) $\centering\begin{split}\mathcal{L}=\mathcal{L}_{SR}+\alpha\mathcal{L}_{Clust}+\beta_{1}\mathcal{L}_{SCL}+\beta_{2}\mathcal{L}_{CKD}+\lambda\mathcal{L}_{Adv}.\end{split}\@add@centering$
where $\alpha$, $\beta_{1}$,$\beta_{2}$ and $\lambda$ are hyper-parameters.
@@ 5.1. Experimental Settings
[…]
@@ 5.2. Overall Performance
Table 2. Overall performance. Bold scores represent the highest results of all methods. Underlined scores stand for the second highest. ”***” denotes the statistical siginificance for $p<0.01$ compared to the best baseline methods with paired $t$-test.
@@
  • $S^{4}$ec achieves remarkable improvement over the strongest baseline ICLRec w.r.t HR@5 by 2.63$\sim$9.26and NDCG@5 by 2.25$\sim$10.13%, respectively. It demonstrates that the proposed framework is dataset agnostic and performs stably given distinct behavior distributions. Further, it is beneficial to employ the adversarial classifier to alleviate the distribution discrepancy between head and tail users in the clustering process.
  • Compared to standard sequential models, $S^{4}$ec indisputably outperforms all benchmarks. Although the Transformer-based encoder achieves the best performance in standard sequential models, it still performs relatively poorly against $S^{4}$ec. Thus, it is critical to utilize self-supervised signals to sufficiently optimize model parameters.
  • Compared to sequential models with SSL, $S^{4}$ec introduces latent prototypes that summarize the semantics of entire user behavior sequences. Although self-supervised signals are utilized in most benchmarks, they focus on augmented behavior-level views to obtain separated user behavior representations but fail to leverage intent-level information from the augmented sequences. As a result, $S^{4}$ec achieves considerable performance gains against this line of research.
  • The performance of $S^{4}$ec leads in each metric compared to ICLRec. The performance of ICLRec is limited as there is no correspondence between two consecutive cluster assignments in ICLRec. Hence, the final prediction layer learned from an assignment may become irrelevant for the following one and thus needs to be reinitialized from scratch at each epoch, which considerably disrupts the model training. On the contrary, $S^{4}$ec maps user behavior sequences to K prototypes in an online fashion. In addition, ICLRec suffers from the head-tail problem, which leads to suboptimal clustering. The results validate the effectiveness of the online clustering and adversarial strategy.
5.3. Head-tail Study
As discussed in Section 1, the uniform distribution is a significant factor in contrastive learning. Thus, we further study the impact of adversarial learning for head-tail on clustering results through the distribution of head rate, defined as the proportion of head sequences within a cluster. As Figure 3 presents, $S^{4}$ec has more compact box plots on all four datasets than ICLRec. This observation indicates the distribution of head sequences across clusters generated by $S^{4}$ec is more uniform than that of ICLRec. When we eliminate the impact of the head-tail problem on clustering, both the performance on head samples and tail samples increases as shown in Figure 4. Therefore, our adversarial strategy optimizes the clustering distribution, leading to better SR performance.
(a) Beauty
(b) Sports
(c) Toys
(d) ML-1M
Figure 3. The quartile boxplots of head rate.
(a) HR@5
(b) NDCG@5
Figure 4. The head/tail performance of $S^{4}$ec compared to ICLRec on the Beauty dataset.
@@ 5.4. Ablation Study
Our proposed $S^{4}$ec contains two essential modules: cluster-aware two-fold self-distillation (CSD) module and Gradient Reversal (GR) layer as an adversarial learning module. To understand the impact of the sub-modules of $S^{4}$ec, we conduct ablation study by removing different sub-module. The results reported in Table 3 are based on experiments conducted in the Amazon Beauty dataset. Similar results are also achieved in other datasets.
  • SR only: standard sequential recommendation model SASRec
  • SR+CSD: apply additional CSD module based on SR
  • SR+CSD+GR: apply both CSD and GR module based on SR and K-means
  • $S^{4}$ec w/o CSD: apply GR module and remove CSD on $S^{4}$ec
  • $S^{4}$ec w/o GR: apply CSD module and remove GR on $S^{4}$ec
  • $S^{4}$ec: the complete configuration $S^{4}$ec
Overall, both the standard SR model and $S^{4}$ec considerably enjoy performance gain from CSD and GR modules. The complete configuration $S^{4}$ec is ahead of SR+CSD+GR by 1.61% in terms of NDCG@20. The main difference between SR+CSD+GR and S4Rec is the employed clustering algorithm. The clustering used by SR+CSD+GR is the traditional k-means algorithm that needs to be executed on the entire dataset, while $S^{4}$ec employs online clustering. That is to say, the recommendation results are not affected by performing clustering within a mini-batch, though the k-means algorithm utilizes all data points. Based on the reported results, we validate the effectiveness of the proposed design choices.
Impact of Cluster-aware Self-distillation. We report that the performance declines significantly as the CSD module is removed, regardless of the sequential recommendation model or the proposed $S^{4}$ec. Concretely, SR without the help of the CSD module suffers a 20.21% performance drop at NDCG@20, and compared to the complete configuration of $S^{4}$ec, the NDCG@20 of $S^{4}$ec without CSD drops 21.39%. We argue that cluster-aware distillation transfers knowledge by extracting additional supervisory signals from the intent-level representation. Therefore, the instance of limited behaviors benefits from intent-level generalization, though the information is limited at the sequence-level. Consequently, the CSD module can efficiently alleviate the problem of insufficient self-supervision signals given users limited behaviors.
Impact of Head-tail Adversarial Learning. The adversarial classifier is indispensable since we aim to prevent the semantics of the sequence length from dominating the clustering process. Through the gradient reversal technique, we report that it brings 2.4% and 5.2% NDCG@20 improvement in terms of SR+CSD+GR and $S^{4}$ec, respectively. Both models enjoy clear resolution on disentangling head-tail semantics, leading to better performance of intent-level clustering.
@@
$S^{4}$ec 0.0519 0.1071 0.0348 0.0505
5.5. Hyper-parameter Sensitivity Study
In this subsection, we investigate and report the impact of a group important hyper-parameters on model performance.
Impact of the number of intent cluster $K$. The number of intent clusters $K$ is vital in the clustering process. We argue that improper choice heavily affects the final performance since $K$ fundamentally influences the user’s intention distribution. In terms of HR@5, $S^{4}$ec enjoys its best performance at $K=128$ and suffers mild loss when $K$ is smaller. On the contrary, the performance drops significantly when $K$ is greater than $128$. We speculate that sufficiently large $K$ results in false negative representations. Concretely, each intent prototype depletes the resolution of the user’s real intentions and further harms the effectiveness of cluster-aware self-distillation.
Impact of multi-task objective weight. As each weight of the objective determines the strength of the gradient during the training process, we vary $\alpha$, $\beta_{1}$,$\beta_{2}$, and $\lambda$ at different scales to analyze the impacts on the performance of $S^{4}$ec. As shown in Figure 5, we observe that performance reaches the peak with the combination $\alpha$=0.01, $\beta_{1}$=0.1,$\beta_{2}$=0.1 and $\lambda$=1.0, and then deteriorates sharply as the weight continues to increase. Coinciding with the discussion of Section 4.3, larger $\lambda$ enforces the CSD module to dispatch the instance into different clusters based on the user’s behavior sequence length. As the boxplot Figure 5(b) demonstrates, head users allocated to each cluster concentrate as $\lambda$ increases.
Table 4. Parameter analysis of the cluster number K.
# of clusters K 32 64 128 256 512
HR@5 0.0510 0.0508 0.0519 0.0481 0.0455
(a) HR@5 on $\alpha$
(b) HR@5 and boxplot on $\lambda$
(c) HR@5 on $\beta_{1}$
(d) HR@5 on $\beta_{2}$
Figure 5. Parameter analysis of $\alpha$, $\lambda$, $\beta_{1}$ and $\beta_{2}$ on the Beauty dataset.

### v2

@@ Abstract.
Sequential recommendation methods play a pivotal role in modern recommendation systems. A key challenge lies in accurately modeling user preferences in the face of data sparsity. To tackle this challenge, recent methods leverage contrastive learning (CL) to derive self-supervision signals by maximizing the mutual information of two augmented views of the original user behavior sequence. Despite their effectiveness, CL-based methods encounter a limitation in fully exploiting self-supervision signals for users with limited behavior data, as users with extensive behaviors naturally offer more information. To address this problem, we introduce a novel learning paradigm, named Online Self-Supervised Self-distillation for Sequential Recommendation ($S^{4}$ec), effectively bridging the gap between self-supervised learning and self-distillation methods. Specifically, we employ online clustering to proficiently group users by their distinct latent intents. Additionally, an adversarial learning strategy is utilized to ensure that the clustering procedure is not affected by the behavior length factor. Subsequently, we employ self-distillation to facilitate the transfer of knowledge from users with extensive behaviors (teachers) to users with limited behaviors (students). Experiments conducted on four real-world datasets validate the effectiveness of the proposed method.
journalyear: 2024copyright: acmlicensedbooktitle: Proceedings of the ACM Web Conference 2024 (WWW ’24), May 13–17, 2024, Singapore, Singaporeisbn: 978-1-4503-XXXX-X/18/06doi: XXXXXXX.XXXXXXXccs: Computing methodologies Learning paradigmsccs: Computing methodologies Neural networksccs: Information systems Recommender systems
@@ 2.1. Self-Supervised Learning for Recommendation
Self-Supervised Learning (SSL) become prevalent in different research areas, including computer vision (Chen and Chu, 2023), natural language processing (Gao et al., 2021), and more. The main target of SSL is to capture high-quality and information-rich representations through the feature itself. There have also been some recent works to apply SSL to sequential recommendations. For example, S3-Rec (Zhou et al., 2020) adopts a pre-training and fine-tuning strategy with four self-supervised tasks, and first proposes to maximize the mutual information between historical items and their attributes. COTREC (Xia et al., 2021) introduces a graph-based recommendation model that utilizes the session-based graph to augment two views, exhibiting the internal and external connectivities of sessions, thereby supervising each other through contrastive learning. While DCN (Lin et al., 2022) broadens its scope by incorporating user sequence data on the item side and construct a dual contrastive learning network to enhance recommendation performance, our paper narrows its focus to user-side modeling. CL4SRec (Xie et al., 2022) introduces three data-level augmentation approaches (crop/mask/reorder, referred to as invasive augmentation methods in the paper) to structure positive views. Later, CoSeRec (Liu et al., 2021a) aims to produce robust augmented sequences based on item denpendencies since random item perturbations may weaken the confidence of positive pairs. ICLRec (Chen et al., 2022) conducts clustering among all user behavior sequences to obtain user’s intent, and optimizes sequential recommendation model by maximizing the mutual information between sequence and corresponding intentions.
@@ 2.2. Intention Learning for Recommendation
Many approaches have been proposed to study users’ intents behind each user’s behavior for improving recommendations (Ma et al., 2020; Chen et al., 2020; Tan et al., 2021). DSSRec (Ma et al., 2020) introduces a sequence2sequence training strategy to capture extra supervision information. An intent variable is employed to extract mutual information between an individual user’s past and future interaction sequences. ICLRec (Chen et al., 2022) learns users’ intent distribution from unlabeled user behavior sequences and optimize SR models with contrastive learning by considering the obtained intents. ISRec (Li et al., 2023) extracts intentions of target user from sequential contexts, then takes complex intent transition into account through the message-passing mechanism on intention graphs.
@@ 4.4. Multi-task Training
We adopt a multi-task strategy where the main next-item prediction, the cluster assignment, the cluster-aware self-distillation and the adversarial learning task are jointly optimized.
@@ 5.1. Experimental Settings
  • Sequential models considering SSL. BERT4Rec  (Sun et al., 2019) proposes a deep bidirectional transformer model to extract both left and right-side behaviors information. S3-Rec  (Zhou et al., 2020) adopts a pre-training and fine-tuning strategy with four self-supervised tasks. CL4SRec  (Xie et al., 2022) introduces three data-level augmentation approaches to construct positive views. This line of works all utilize the transformer as sequence encoder but adopt distinct constrastive learning tasks.
@@ 5.2. Overall Performance
Table 2. Overall performance. Bold scores represent the highest results of all methods. Underlined scores stand for the second highest. ”*” denotes the statistical siginificance for $p<0.01$ compared to the best baseline methods with paired $t$-test.
@@ 5.3. Ablation Study
Our proposed method contains two essential modules: cluster-aware two-fold self-distillation (CSD) module and Gradient Reversal (GR) layer as an adversarial learning module. To understand the impact of the sub-modules of $S^{4}$ec, we conduct ablation study by removing different sub-module. The results reported in Table 3 are based on experiments conducted in the Amazon Beauty dataset. Similar results are also achieved in other datasets.